OpenStudy (anonymous):

Keonna flipped a coin 3 times. What is the probability that she flipped heads all 3 times? A.1/2 B.1/8 C.1/6 D.3/8

OpenStudy (anonymous):

B 1/2 * 1/2 * 1/2 = 1/8

OpenStudy (cwrw238):

these probabilities are independent so its 1/2 * 1/2 * 1/2 = 1/8